We consider a Standard Brownian Motion W={Wt,t>=o}, show that for s<t, Ws|Wt=x the conditional distribution of the process given a future valueWt=x

We consider a standard Brownian motion W W,t20) Show that for s < t, W /Wt-x the conditional distribution of the process given a future value Wi is given by the following Normal distribution:
Answer #1

If X and Y are jointly normal random variables with parameters mu_x, sigma_x^2, mu_y, sigma_y^2, ho , then Y | X is normally distributed with

E[Y | X = x] = mu_y + ho sigma_y (x - mu_x)/sigma_x

Var[Y | X= x] = (1 - ho^2) sigma_y^2

Let X = Wt and Y = Ws , so we have X ~ N(0, t) and Y ~ N(0, s) and

Cov(X, Y) min(t, s s Vtys (s < t)


ElyIX =x] = 0 +ys.VS((x _ 0)/Vt) =.t

Var[X | Y = y] = left ( 1 - rac{s}{t} ight )* s = rac{s}{t} left ( t - s ight )

Thus, the distribution of Ws | Wt = x is,

W_s | W_t = x sim N left ( rac{s}{t}x, rac{s}{t} left ( t - s ight ) ight )